OpenStudy (anonymous):

how to get sum of 3/4 + 1/6

myininaya (myininaya):

find a common denominator what is the least common multiple of 4 and 6

myininaya (myininaya):

4*3=12 6*2=12

OpenStudy (amistre64):

add them together :)

OpenStudy (anonymous):

12

myininaya (myininaya):

gj rem

OpenStudy (anonymous):

so do i multiply the top numbers too

myininaya (myininaya):

3/4=9/12 1/6=2/12 right rem

myininaya (myininaya):

what do you get when you add those two fractions?

OpenStudy (anonymous):

11/12

myininaya (myininaya):

gj task completed!

OpenStudy (anonymous):

thank you